The half-life of strontium-91, 91


38Sr is 9.70 h. Find


(a) its decay constant, and (5pts)

(b) for an initial 1.00-g sample, the activity after 15 hours.


1
Expert's answer
2021-10-30T01:15:05-0400

Solution:

(a):

The time required for half of the original population of radioactive atoms to decay is called the half-life.

The relationship between the half-life (t1/2) and the decay constant (λ) is given by t1/2 = 0.693 / λ.

Hence,

λ = 0.693 / t1/2 

λ = (0.693) / (9.70 h) = 0.0714 h-1

λ = 0.0714 h-1


(b):

Calculate the number of atoms in 1.00 g of strontium-91:

No = m × Na / M

No = (1.00 g × 6.022×1023) / (91 g/mol) = 6.618×1021

Now, the activity of the sample:

A = No × λ × e–λt,

where:

A = ending activity

No = starting number of atoms

e = constant = 2.71828

λ = decay constant

t = elapsed time


Hence,

A = (6.618×1021) × (0.0714 h-1) × (2.71828)–0.0714 × 15 = 1.62×1020 Bq

A = 1.62×1020 Bq


Answers:

(a): λ = 0.0714 h-1

(b): A = 1.62×1020 Bq